OpenStudy (anonymous):

What is the sum of the solutions of ? A. 1 1/2 B. -2 C. 2 D. 0

OpenStudy (anonymous):

\[2|x-1|-4= -2\]

OpenStudy (shubhamsrg):

any attempt ?

OpenStudy (anonymous):

pick d but got it wrong

OpenStudy (shubhamsrg):

you just guessed d) ??

OpenStudy (anonymous):

no

OpenStudy (shubhamsrg):

so what was your logic behind saying "d" ?

Directrix (directrix):

2|x-1|-4= -2 2|x-1| = -2 + 4 2|x-1| = 2 |x-1| = 1 So, (x-1) = 1 OR (x-1) = -1 Solve each equation above and add the solutions. @jbarber